OpenStudy (anonymous):

use the method of substitution to solve the system of linear equations. 2x+4=2/3 y=3x-1 The solution is ?

OpenStudy (anonymous):

solve the top equation for x. plug the answer for x into the bottom equation and post your answer for y. I'll check it.

OpenStudy (anonymous):

Im not sure how to do it

OpenStudy (anonymous):

2x+4 = 2/3 \[2x = \frac{ 2 }{3}-4\] \[2x = -3\frac{ 1 }{3 }\] \[2x = \frac{ 10 }{3}\] \[x = \frac{ 10 }{ 6 }\] \[x = \frac{ 5 }{ 3 }\] now you can solve the lower part